# CORAL: Learning Consistent Representations across Multi-step Training with Lighter Speculative Drafter

### v2

@@ Abstract
Yepeng Weng, Dianwen Mei, Huishi Qiu Xujie Chen, Li Liu, Jiang Tian, Zhongchao Shi Lenovo Research
@@ 3.3 LM Head Router
It is evident that, although the LM head router reduces the latency of the draft model, it comes at the cost of a slight decrease in acceptance length $\tau$ due to imperfect routing accuracy. Based on Equation (5) and (6), the LM head router gets its best performance when 1) the LM head accounts for a significant portion of the latency of draft model 2) the latency ratio between the draft model and the target model is substantial. Therefore, we only apply the LM head router to models with large vocabularies (Qwen2.5, Llama3) and relatively small sizes (7B, 14B).
We adopt a two-stage training strategy, where we first train the draft model following the standard training procedure (either single-step or multi-step), and then fix the weights of draft model and train the router separately. For further discussion, please refer to Appendix D.
@@ 4.1 Experimental Setup
Tasks. We choose multiple datasets covering three tasks, including MT-BenchZheng et al. (2023) for multi-turn dialogue, GSM8KCobbe et al. (2021) for mathematical reasoning, and HumanEvalChen et al. (2021) for code generation. For 7B/14B models, experiments are conducted with batch size of 1 on a single NVIDIA A6000 48G GPU. For Llama3-70B, we use 4$\times$×A6000 GPUs due to memory requirements.
@@
L2-13B EAGLE 3.93/3.04$\times$ 3.23/2.35$\times$ 4.51/3.47$\times$ 3.47/2.56$\times$ 4.01/3.10$\times$ 3.51/2.59$\times$ 4.15/3.20$\times$
@@
L2-7B EAGLE 3.80/2.67$\times$ 3.21/2.10$\times$ 4.29/3.04$\times$ 3.55/2.33$\times$ 3.84/2.73$\times$ 3.48/2.30$\times$ 3.87/2.81$\times$
@@
L3-70B EAGLE 2.87/2.24$\times$ 2.62/2.02$\times$ 3.73/2.93$\times$ 3.45/2.67$\times$ 3.46/2.71$\times$ 3.23/2.50$\times$ 3.35/2.63$\times$
@@
L3-8B EAGLE 2.63/1.65$\times$ 2.30/1.35$\times$ 3.65/2.29$\times$ 3.13/1.85$\times$ 3.47/2.18$\times$ 3.05/1.78$\times$ 3.25/2.04$\times$
@@
Q2.5-14B EAGLE 2.63/1.83$\times$ 2.33/1.55$\times$ 3.31/2.31$\times$ 2.82/1.88$\times$ 3.62/2.52$\times$ 3.21/2.16$\times$ 3.19/2.22$\times$
@@
Q2.5-7B EAGLE 2.53/1.56$\times$ 2.17/1.23$\times$ 3.04/1.87$\times$ 2.62/1.49$\times$ 3.32/2.05$\times$ 2.86/1.63$\times$ 2.96/1.83$\times$
@@ 4.2.2 Ablation Study on CSRA
We adjust the number of training steps and make a more detailed comparison with HASS. Since CSRA and HASS employ the same draft model, the inference overheads are identical, we therefore compare the acceptance length only. The results in Table 3 show that CSRA consistently outperforms HASS under different training steps.
MT-bench HumanEval GSM8K
step HASS CSRA HASS CSRA HASS CSRA
2 4.41 4.53 5.24 5.35 4.50 4.60
3 4.48 4.57 5.31 5.43 4.56 4.70
4 4.46 4.58 5.39 5.55 4.58 4.70
Table 3: Acceptance length of Llama3-8B under different alignment steps. Step-3 is the default setting.
To provide a more intuitive measure of the alignment between the draft model and the target model, we compare the acceptance rates $\alpha$ of HASS and CSRA at different timesteps during inference, as shown in Figure 6. The results show that CSRA generally outperforms HASS at different timesteps.
@@ 5 Related Work
There has been a significant amount of work in accelerating LLMs. Some methods focus on reducing the number of parameters, such as low-bit quantization Dettmers et al. (2022); Frantar et al. (2023); Xiao et al. (2023); Lin et al. (2024), and model distillation Gu et al. (2024); Ko et al. (2024); Zhong et al. (2024). Recently, some studies have also explored activating only a subset of model parameters during inference to reduce memory access cost Du et al. (2022); Fedus et al. (2022). Speculative decoding Chen et al. (2023); Leviathan et al. (2023) leverages the memory-bound nature of decoder-only LLMs and achieves lossless acceleration using a drafting-verification framework.
@@ 6 Conclusion
This paper proposes CORAL, an efficient speculative decoding method. We introduce Cross-Step Representation Alignment, which effectively mitigates training-inference misalignment and improves the accuracy of speculation. Additionally, we propose the LM head router, a plug-and-play module designed to reduce the latency of the draft model. We compare CORAL with other state-of-the-art methods on various LLMs and datasets, and the results show that CORAL achieves the best speedup performance.
@@ Appendix B Training Details
We utilize a fixed dataset of 68,000 examples from ShareGPT [^3] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna
@@ Appendix C Single-step Training with CSRA
Appendix D Discussion on LM Head Router
In this section, we will discuss some issues of LM head router.
@@ Appendix E Discussion on System Prompt
Appendix F Licenses of Artifacts
@@
Appendix G Use of AI Assistants
We use Llama3.2-90B to assist with grammar checks and text polishing in the writing of this paper.

### v3

@@ Abstract
Yepeng Weng, Dianwen Mei, Huishi Qiu Xujie Chen, Li Liu, Jiang Tian, Zhongchao Shi AI Lab, Lenovo Research {wengyp1, meidw1, qiuhs1, chenxj23, liuli16, tianjiang1, shizc2}@lenovo.com
@@ 3.3 LM Head Router
It is evident that, although the LM head router reduces the latency of the draft model, it comes at the cost of a slight decrease in acceptance length $\tau$ due to imperfect routing accuracy. Based on Equations (5) and (6), the LM head router gets its best performance when 1) the LM head accounts for a significant portion of the latency of draft model 2) the latency ratio between the draft model and the target model is substantial. Therefore, we only apply the LM head router to models with large vocabularies (Qwen2.5, Llama3) and relatively small sizes (7B, 14B).
We adopt a two-stage training strategy, where we first train the draft model following the standard training procedure (either single-step or multi-step), and then fix the weights of draft model and train the router separately. For further discussion, please refer to Appendix F.
@@ 4.1 Experimental Setup
Tasks. We choose multiple datasets covering three tasks, including MT-BenchZheng et al. (2023) for multi-turn dialogue, GSM8KCobbe et al. (2021) for mathematical reasoning, and HumanEvalChen et al. (2021) for code generation. For 7B/14B models, experiments are conducted with batch size of 1 on a single NVIDIA A6000 48G GPU. For Llama3-70B-Instruct, we use 4$\times$×A6000 GPUs due to memory requirements.
@@
L2-13B EAGLE 3.93/3.04$\times$ 3.64/2.65$\times$ 4.51/3.47$\times$ 4.24/3.13$\times$ 4.01/3.10$\times$ 3.84/2.83$\times$ 4.15/3.20$\times$
@@
L2-7B EAGLE 3.80/2.67$\times$ 3.62/2.37$\times$ 4.29/3.04$\times$ 3.96/2.60$\times$ 3.84/2.73$\times$ 3.74/2.48$\times$ 3.87/2.81$\times$
@@
L3-70B EAGLE 2.87/2.24$\times$ 2.67/2.06$\times$ 3.73/2.93$\times$ 3.53/2.74$\times$ 3.46/2.71$\times$ 3.26/2.52$\times$ 3.35/2.63$\times$
@@
L3-8B EAGLE 2.63/1.65$\times$ 2.40/1.41$\times$ 3.65/2.29$\times$ 3.29/1.92$\times$ 3.47/2.18$\times$ 3.22/1.89$\times$ 3.25/2.04$\times$
@@
Q2.5-14B EAGLE 2.63/1.83$\times$ 2.44/1.62$\times$ 3.31/2.31$\times$ 3.12/2.10$\times$ 3.62/2.52$\times$ 3.46/2.33$\times$ 3.19/2.22$\times$
@@
Q2.5-7B EAGLE 2.53/1.56$\times$ 2.25/1.27$\times$ 3.04/1.87$\times$ 2.79/1.58$\times$ 3.32/2.05$\times$ 3.00/1.72$\times$ 2.96/1.83$\times$
@@ 4.2.1 Effectiveness
MT-bench HumanEval GSM8K Average
step HASS CSRA HASS CSRA HASS CSRA HASS CSRA
2 4.41/2.41$\times$ 4.53/2.48$\times$ 5.24/2.86$\times$ 5.35/2.90$\times$ 4.50/2.47$\times$ 4.60/2.52$\times$ 4.72/2.58$\times$ 4.83/2.63$\times$
3 4.48/2.45$\times$ 4.57/2.50$\times$ 5.31/2.89$\times$ 5.43/2.95$\times$ 4.56/2.51$\times$ 4.70/2.58$\times$ 4.78/2.62$\times$ 4.90/2.68$\times$
4 4.46/2.44$\times$ 4.58/2.51$\times$ 5.39/2.93$\times$ 5.55/3.00$\times$ 4.58/2.54$\times$ 4.70/2.57$\times$ 4.81/2.64$\times$ 4.94/2.69$\times$
Table 3: Acceptance length and speedup ratio of Llama3-8B under different alignment steps.
@@ 4.2.2 Ablation Study on CSRA
We conduct a more detailed comparative analysis with HASS under different training steps. According to HASS, further increases in the number of training steps (i.e., training steps $\geq$ 5) do not necessarily lead to improvements in acceptance length. Therefore, we focus our comparison on the cases where the number of training steps is set to 2, 3 (default), and 4.
The results summarized in Table 3 demonstrate that CSRA consistently outperforms HASS with the same number of training steps. Furthermore, from the perspective of training cost, CSRA achieves a performance comparable to HASS (step=4) using only 2 training steps, demonstrating that CSRA offers a substantial advantage in terms of training efficiency. We also compare the acceptance rates $\alpha$ of HASS and CSRA at different timesteps during inference, as shown in Figure 6. The results show that CSRA generally outperforms HASS at different timesteps.
@@ 4.2.3 Ablation Study on LM Head Router
Let us discuss the effectiveness of the LM head router from another aspect using Llama3-8B. Since the number of activated LM head groups may vary due to tree decoding, we can estimate the latency of the draft model based on the ratio between acceptance length and speedup. Specifically, according to Table 1, the latency of the draft model is approximately 10% that of the target model. Therefore, ideally (assuming $L_{t}^{\prime}=L_{t}$ in Equation 1), the ratio between acceptance length and speedup should be 1.6, which means that within one drafting-verification cycle, the target model is invoked once while the draft model runs six times. However, in practical scenarios, the latency of the target model increases by approximately 19% (from 26ms to 31ms) when inferring 60 tokens in parallel compared to generating a single token. Taking into account this factor, the actual ratio between acceptance length and speedup will increase to approximately 1.8.
Our experiments also confirm this estimation. Without the LM head router, the ratio between acceptance length and speedup is approximately $4.90/2.68\approx 1.83$. In contrast, when the LM head router is adopted, this ratio decreases to $4.63/2.87\approx 1.61$. This indicates that the average latency of the draft model is only $(0.6-0.22)/0.6\approx 63\%$ of its original latency, demonstrating the efficacy of the LM head router.
@@ 5 Related Work
There has been a significant amount of work in accelerating LLMs. Some methods focus on reducing the number of parameters or memory access, such as low-bit quantization Dettmers et al. (2022); Frantar et al. (2023); Xiao et al. (2023); Lin et al. (2024), and model distillation Gu et al. (2024); Ko et al. (2024); Zhong et al. (2024). Recently, some studies have also explored activating only a subset of model parameters during inference to reduce memory access cost Du et al. (2022); Fedus et al. (2022). Speculative decoding Chen et al. (2023); Leviathan et al. (2023) leverages the memory-bound nature of decoder-only LLMs and achieves lossless acceleration using a drafting-verification framework.
@@ 6 Conclusion
This paper proposes CORAL, an efficient speculative decoding method. We introduce Cross-Step Representation Alignment, which effectively mitigates training-inference misalignment and improves the accuracy of speculation. Additionally, we propose the LM head router, a plug-and-play module designed to reduce the latency of the draft model. We compare CORAL with other state-of-the-art methods on various LLMs and datasets, and the results show that CORAL surpasses existing methods, such as EAGLE-2 and HASS, demonstrating the effectiveness of our method.
@@
Acknowledgments
We would like to thank Lenovo Model Factory team for providing computing resources. Special thanks to Xiaoyue Mi from the Institute of Computing Technology, Chinese Academy of Sciences, Penghui Yang from Nanyang Technological University, and Henry Zheng from Tsinghua University for their valuable suggestions during the writing of this paper.
@@ Appendix B Training Details
We use a fixed dataset of 68,000 examples from ShareGPT [^3] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna
@@ Appendix C Single-step Training with CSRA
Appendix D Discussion on the discrepancies between different training steps
To better illustrate the discrepancies between representations from multiple training steps, we measure the InfoNCE between features from different steps. Please note that absolute distance metrics (such as L1 or cosine distance) are not ideal measurements, as absolute distances fail to represent the distinguishability between different features. In contrast, InfoNCE transforms cosine similarity into a probability distribution, effectively reflecting the relative distances between features, which is more crucial for prediction accuracy. Therefore, InfoNCE serves as a more appropriate metric.
EAGLE-2 step 0 step 1 step 2 step 3
step 0 - - - -
step 1 1.5668 - - -
step 2 1.8415 1.4843 - -
step 3 2.0391 1.6559 1.4876 -
HASS step 0 step 1 step 2 step 3
step 0 - - - -
step 1 1.4577 - - -
step 2 1.6117 1.3816 - -
step 3 1.7290 1.4876 1.3924 -
CSRA step 0 step 1 step 2 step 3
step 0 - - - -
step 1 0.9545 - - -
step 2 1.1044 0.8395 - -
step 3 1.2179 0.9381 0.8173 -
Table 6: InfoNCE between features from different training steps. The temperature is set to 0.07, which is aligned with our setting in CSRA training.
We randomly select 100 samples from ShareGPT and evaluate the differences in output features across 4 steps. The following are the InfoNCE between features of different steps for EAGLE-2, HASS, and CSRA. For instance, the first column represents the InfoNCE between the output features of step 1 and step 2 to 4. Clearly, the differences in features between steps increase gradually as the number of steps grows. Since HASS employs multi-step training, the differences between steps are smaller compared to EAGLE-2. Moreover, our method significantly reduces the discrepancies between different steps, achieves higher similarity between positive features and enhances the discriminative power of negative features, ensuring relatively consistent performance across all steps during inference.
@@ Appendix E Discussion on System Prompt
Appendix F Discussion on LM Head Router
In this section, we will discuss some issues of LM head router.
Alpaca Natural Q. CNN/DM
model method $\tau$ SR $\tau$ SR $\tau$ SR
L2-7B EAGLE-2 4.51 2.88$\times$ 4.10 2.61$\times$ 4.12 2.40$\times$
HASS 4.87 3.11$\times$ 4.41 2.80$\times$ 4.44 2.57$\times$
CORAL 4.96 3.15$\times$ 4.44 2.84$\times$ 4.54 2.62$\times$
L3-8B EAGLE-2 4.33 2.39$\times$ 3.37 1.86$\times$ 3.82 1.98$\times$
HASS 4.77 2.56$\times$ 3.59 1.98$\times$ 4.06 2.16$\times$
CORAL 4.79 2.58$\times$ 3.63 2.00$\times$ 4.16 2.20$\times$
CORAL w/ r. 4.49 2.74$\times$ 3.28 2.06$\times$ 3.61 2.16$\times$
Q2.5-7B EAGLE-2 3.93 2.17$\times$ 3.13 1.73$\times$ 3.33 1.78$\times$
HASS 4.19 2.31$\times$ 3.30 1.82$\times$ 3.57 1.90$\times$
CORAL 4.29 2.35$\times$ 3.38 1.86$\times$ 3.72 1.97$\times$
CORAL w/ r. 4.11 2.60$\times$ 3.15 1.99$\times$ 3.28 1.99$\times$
Table 8: Additional results on Alpaca, Natural Questions and CNN/DM dataset. We provide the results on Llama2-7B-chat, Llama3-8B-Instruct and Qwen2.5-7B-Instruct. The temperature is set to 0.
nchmark.html
Small vocabulary and super large LLMs. Let’s take Llama3-70B on MT-bench as an example. In our experiments on CORAL, the time consumption of the target/draft model is 4105s/561s, meaning that the draft model accounts for only 12% of the entire drafting-verification cycle (for Llama3-8B, this figure is approximately 33%). Although the LM head of the draft model still constitutes a significant portion of the drafting latency, its overall contribution to the entire cycle is only 5–6% (while for Llama3-8B, it is nearly 20%). If a router is used, the time consumption of the target/draft model becomes 4477s/440s, resulting in only a 3% reduction in the entire cycle. However, the acceptance length decreases from 4.23 to 3.93, a drop of 9.3%, and the speedup decreases from 2.79$\times$ to 2.69$\times$.
A similar conclusion applies to Llama2-7B. Since the latency of the LM head does not constitute a large part of the total latency, using a router on Llama2 is not a good choice.
Appendix G Additional Experiments
Here we present some additional experimental results on Alpaca Taori et al. (2023), Natural Questions Kwiatkowski et al. (2019) and CNN/DM Nallapati et al. (2016) datasets.
Appendix H Licenses of Artifacts